OpenStudy (anonymous):

convert the scale to a unit rate. show your work. interpret the meaning of the unit rate. 3 7 - - 10 in. = 8 mi ?

OpenStudy (therealmeeeee):

ok ik tthis

OpenStudy (therealmeeeee):

3/10 in.= 7/8 mi you can cross multiply them what do you get?

OpenStudy (therealmeeeee):

Multiply both sides by 10/3

OpenStudy (therealmeeeee):

3/10 = 7/8 1 - 10/10)= X 3/10x = 7/8 x= 7/8 divided by 3/10 = 7/8 x 10/3 = 70/24 = 2.916 1 unit = 2.916 units

OpenStudy (therealmeeeee):

there ya go di u get it now??????????

OpenStudy (anonymous):

yeah thank you very much~! you were very helpful

OpenStudy (therealmeeeee):

my medal and fan

OpenStudy (therealmeeeee):

u have to give me a medal

OpenStudy (therealmeeeee):

my medal